a hockey skater turns in a circular path with a radius of 31.0 m while going 10.5 m/s. If he experiences a centripetal force of 565 N, what is the skater’s mass?


1
Expert's answer
2021-03-02T18:04:59-0500

From the second Newton's law, the mass can be expressed as follows:


"m = \dfrac{F}{a}"

where "F = 565N" is the net force acting on the skater, "a" is they acceleration. In turn, the centripetal acceleration is given as follows:


"a = \dfrac{v^2}{R}"

where "v = 10.5m\/s" is the speed, and "R = 31m" is the radius. Thus, obtain:


"m = \dfrac{FR}{v^2}\\
m = \dfrac{565\cdot 31}{10.5^2}\approx 158.9kg"

Answer. 158.9 kg.